OpenStudy (anonymous):

For the polynomial ƒ(x)=1-2x+5x^4 as x-->∞,ƒ(x)-->-∞ true or false

OpenStudy (anonymous):

Pick a really big number

OpenStudy (anonymous):

plug it in for x.

OpenStudy (anonymous):

Or you can just think about what would happen if x gets really big and positive.

OpenStudy (anonymous):

1 isn't gonna change as x changes. -2x will get big and negative as x gets big and positive +5x^4 will get big and positive as x gets big and positive. So will -2x get more negative than 5x^4 gets positive?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Does the graph y=5x^4 go up faster than y=2x

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Right, so 5x^4 is going to go up faster than -2x is going to go down. Therefore the result will be +\(\infty\) not \(-\infty\)

OpenStudy (anonymous):

I asuume this is a calc question...\[\lim_{x \rightarrow \infty} (1-2x+5x ^{4}) = \lim_{x \rightarrow \infty} f'(x)= \lim_{x \rightarrow \infty} (-2+20x ^{3}) =+\infty\]